Because ofthe drought, Country Bank Ltd has experienced large losses on its rural loan portfolio and is unable to payits next three annual interest payments on its recent issue of unsecured notes. The notes are of $10,000 face value each, mature in January, 2024 and bear a yearly interest coupon payment of 9%. The Bank paid the interest due this month (January, 2019), and following a meeting of creditors, arranged to defer payment of the next three interest coupons due in January 2020, January, 2021 and January, 2022 respectively. Under the arrangement with creditors, the bank will pay the remaining interest coupons (due in January, 2023 and January, 2024) on their due dates, and paythe three deferred coupons (without interest) along with the normal final interest payment and face value of the notes on the maturity date. Country Bank Ltds notes are now seen as nsky, and require an 18% per annum retum. REQUIRED Calculate the current value of each unsecured note issued by Country Bank Ltd.
Answer #1
A B C D E F G
1 Years Jan 2020 Jan 2021 Jan 2022 Jan 2023 Jan 2024
2 Coupon interest $900 $900 $900 $900 $900
3 Coupon payment $900 $900
4 Deferred payment $2700
5 Principal repayment $10000
6 Cash payment $0 $0 $0 $900 $13600
7 PVF @ 18% 0.847 0.718 0.608 0.515 0.437
8 PV of cash flow $0 $0 $0 $463.5 $5943.2
9 Current Value $6406.7
10
11 Refer below formulas used above:
12 Years Jan 2020 Jan 2021 Jan 2022 Jan 2023 Jan 2024
13 Coupon interest =10000*9% =B2 =C2 =D2 =E2
14 Coupon payment =D2 =E2 =F2
15 Deferred payment =B2+C2+D2
16 Principal repayment
17 Cash payment =SUM(B3:B5) =SUM(C3:C5) =SUM(D3:D5) =SUM(E3:E5) =SUM(F3:F5)
18 PVF @ 18% =1/(1.18)^1 =1/(1.18)^2 =1/(1.18)^3 =1/(1.18)^4 =1/(1.18)^15
19 PV of cash flow =B6*B7 =C6*C7 =D6*D7 =E6*E7 =F6*F7
20 Current Value =SUM(B8:F8)